# Federated Graph Learning for Cross-Domain Recommendation

### v1

@@ 1 Introduction
In this paper, we focus on a more generic scenario of Broader-Source Cross-Domain Recommendation (BS-CDR), which integrates knowledge from more than two source domains while preserving privacy. Given the diverse nature of user preferences, it is essential to gain a more holistic understanding of user interests by incorporating user behaviors from diversified domains  [10; 11]. For example, in Figure 1(a), a user who enjoys certain types of books might also enjoy movies, toys, and games in similar genres. However, incorporating more domains while preserving privacy poses challenges to counteract negative transfer (NT), which is a phenomenon of transferring knowlege from a source domain that negatively impacts the recommender model performance in the target domain [12]. Suppose the Books domain in Figure 1(a) is the target domain. The Clothing domain is causing NT, because of the domain discrepancy. While the Music domain is supposed to transfer positive knowledge, it might also lead to NT because of lossy privacy-preserving techniques applied to broader source domains. As a result, the influx of negative knowledge accumulated from source domains will poison the model performance of the target domain in BS-CDR scenarios.
To mitigate the NT issue, attention mechanisms have been widely leveraged, either in an explicit (e.g., determine domain attentions by predefined domain features [13; 14]) or implicit manner (e.g., employ hyper-parameters [7; 15]). Another work [16] ensures positive transfer by passing only user-related features. However, existing methods cannot be directly applied to BS-CDR due to two major challenges. First, inadequate privacy preservation (CH1). Both intra-domain and inter-domain privacy must be carefully considered in BS-CDR. As depicted in Figure 1(a), BS-CDR relies on extensive knowledge transfer, risking simultaneous privacy leakages across broader source domains (inter-domain privacy) [9; 17; 18; 19]. Additionally, concerns over centralized data storage may prevent users from sharing sensitive rating data (intra-domain privacy). Second, accumulative negative transfer (CH2). Adjusting hyper-parameters for a large number of source domains in BS-CDR scenarios is extremely difficult, as well as predefined domain features cannot accommodate complex domain diversities. For instance, the impact of NT can inevitably intensify with an increasing number of source domains (Figure 1(b)) [2]. In addition, the use of various lossy privacy-preserving techniques can further degrade the quality of transferred knowledge, complicating the achievement of positive transfer. Consequently, the performance of CDR models can decline to levels lower than those of single-domain models, as shown in Figure 1(b).
@@
Figure 1: (a) In order to obtain accurate recommendations in the Books domain, we aim to exploit user preferences (i.e., knowledge of foreign domains should be fully utilized, e.g. Movie, Toys, and Games domains). However, with the influence of lossy privacy-preserving techniques, the results of the transfer could be negative (e.g., the Music domain with low-quality data). (b) There is a diminishing marginal effect on the growth rate of the model performance with pure positive knowledge, while NT accumulates with an increasing number of source domains. Consequently, the performance of existing methods declines and is worse than that of a single domain model.
@@
  • We propose two key model, i.e., the positive knowledge transfer module and the positive knowledge activation module, to address NT and privacy preservation simultaneously in BS-CDR;
  • We conduct extensive experiments on the Amazon datasets that confirm the effectiveness of FedGCDR in terms of recommendation accuracy.
@@ 2.2 Federated recommendation
Recently, FL [27; 28; 29; 30; 31] has been widely adopted to tackle the privacy issue in recommender system. Chai et al. [32] adopts FL to classic matrix factorization algorithm and utilize homomorphic encryption to avoid the potential threat of privacy disclosure. Later, Wu et al. [33] explores the application of federated graph neural networks (GNN) models to improve the recommendation quality and ensure user privacy. To utilize sensitive social information, Liu et al. [8] adopts local differential privacy (LDP) and negative sampling. More recent studies use VFL to protect company’s privacy in recommender system. Mai et al. [34] utilizes random projection and ternary quantization to ensure privacy preservation in VFL. In CDR, Chen et al. [9] designs a dual-target VFL CDR model with orthogonal mapping matrix and LDP for organizations’ privacy preservation. Liu et al. [35] designs a graph convolutional networks (GCN)-based federated framework to learn user preference distributions for more accurate recommendations. To ensure user privacy in CDR, Liu et al. [6] utilizes a VAE-based federated model to mine user preference with data stored locally. Wu et al. [7] designs a personal module and a transfer module to provide personalised recommendation while preserving user privacy. These existing works, especially federated CDR frameworks, only protect the privacy of either individual users (intra-domain privacy) or the organizations (inter-domain privacy) but not both at the same time. We aim to provide both intra-domain and inter-domain privacy.
@@ 3.2.1 Overview
The overall framework of FedGCDR is shown in Figure 2. FedGCDR follows a Horizontal-Vertical-Horizontal (HVH) pipeline and its two horizontal stages ensure the intra-domain privacy (i.e., individual user level) privacy. Our two key modules focus on the vertical stage and the second horizontal stage: (1) The positive knowledge transfer module preserves the inter-domain privacy by DP and alleviates NT by feature mapping. (2) The positive knowledge activation module filters out potential harmful or conflicting knowledge from the source domains. Specifically, we expand the local graph of the target domain by virtual social links, such that the target domain GAT model could generate reliable domain attention based on the expanded graph. After target domain GAT model training, we further mitigate NT by adopting a fine-tuning stage.
@@
The HVH pipeline contains three stages with switching federated settings. The first horizontal stage refers to the source domain training in which source domain servers individually interacts with its domain users (clients). The private rating matrices are stored within each client, while the clients exchange model and gradients to train a domain-specific global model. The next two stages correspond to our two key modules (vertical positive knowledge transfer module and horizontal positive knowledge activation module), which we will cover in detail in the following subsections. It’s important to note that the vertical positive knowledge transfer module is completely computed in each client’s user space (their personal devices). This is because the needed source domain knowledge can be extracted from local source models on each client which are distributed during the first horizontal source domain training stage.
Following the HVH pipeline, we achieve: (1) Privacy enhancement. The two horizontal stages can provide intra-domain privacy preservation, while we further ensure inter-domain privacy by applying privacy technism to the vertical stage. In the mean time, servers are not involved in the knowledge transfer process (i.e., the positive knowledge transfer module), making them unaware of user interactions in other source domains. (2) Communication efficiency. Cross-domain knowledge transfer does not require additional communication overhead.
@@
We adopt a GAT-based [36; 37; 38] federated framework as the underlying model for our intra-domain recommender system. The horizontal paradigm avoids centralized storage of user ratings to ensure intra-domain privacy (CH1). In the initial step, each user and item is offered an ID embedding of size $d$, denoted by $\mathbb{e}^{0}_{u}$, $\mathbb{e}^{0}_{v}\in\mathbb{R}^{d}$ respectively. The embedding is passed through $L$ message propagation layers [39; 40; 41]. For the $l$-th layer:
@@
In each source domain, the domain server and corresponding users collaboratively train a GAT-based federated model. The training process follows the horizontal federated learning (HFL) paradigm in which only the model and gradients are exchanged considering intra-domain privacy. We will not detail the horizontal federation model (e.g., further privacy guarantee and more high-order information) as it is a well established FL model and not our novel contribution. This model can be replaced by other GAT-based FL models [33; 43] as well.
@@ 3.2.2 Positive knowledge transfer module
After the source domain training, we obtain a series of source models in individual client’s user space. Our positive knowledge transfer module then prepares positive knowledge to be transferred from each source domains ${\mathcal{D}}^{S}$ to the target domain $\mathcal{D}^{T}$, while protecting inter-domain privacy (CH1). Specifically, suppose a individual user (client) $u$ and a source domain ${\mathcal{D}}^{S_{i}}$, we transfer the user $u$’s embedding matrix $\mathbb{X}_{S_{i}}$ in $\mathbb{R}^{L\times d}$ Take the row $l$ of the matrix (i.e., $\mathbb{x}^{l}_{S_{i}}$) as an example, it is the user $u$’s embedding output by the $l$-th message propagation layer. In an ideal scenario (i.e., we transfer totally positive knowledge without taking inter-domain privacy into account) [6], embedding matrices from different source domains can be directly used to enhance target domain local training in client $u$. By utilizing the source domain embeddings, $u$’s final target domain embedding $\mathbb{e}_{T}^{l}$ of layer $l$ is:
@@
where $f_{T}(\cdot)$ is the function that the target domain aggregates the knowledge of the source domains and we will give its expression in Subsection 3.2.3. In this process, the transfer of knowledge between domains takes place entirely in the user $u$’s local space. Such a fully localized mode of knowledge transfer avoiding the additional communication overhead and potential privacy issues [6]. However, this direct emebddings transfer does not meet the privacy and NT constrains in BS-CDR scenarios.
@@
In existing CDR frameworks, the user or item embedding was shared as knowledge [9; 15; 6], which neglects inter-domain privacy. In a GNN-based approach, such direct transfers are subject to privacy attacks. Each message propagation layer can be viewed as a function with user and item embeddings as input. An attacker can easily obtain the user’s private rating matrix based on these embeddings. We apply DP to the souce domain embeddings $\mathbb{x}_{S_{i}}$ [20; 44] to safeguard inter-domain privacy.
@@
User features (i.e., user embeddings) could represent personal preferences and are influenced by domain features (e.g., items). The discrepancy of domains leads to the heterogeneity of feature space between domains which means that source domain embeddings cannot be utilized directly by the target domain. Man et al. [15] show that there exists an underlying mapping relationship between the latent user matrix of different domains, which can be captured by a mapping function. In order to alleviate NT, we adopt a series of MLP to explore mapping functions for each source domain. Adding Gaussian noise and feature mapping, Equation (5) becomes:
@@ 3.2.3 Positive knowledge activation module
After the aforementioned operations, the target domain obtains a list of source domain matrices $\hat{\mathbb{X}}_{S_{1}},\hat{\mathbb{X}}_{S_{2}},...,\hat{\mathbb{X}}_{S_{M-1}}$. The row of the matrices represent $MLP_{i}(\hat{\mathbb{x}}_{S_{i}}^{l})$. It is worth noting that for source domains where a user has no rating, $\hat{\mathbb{X}}_{S_{i}}$ is a Gaussian noise matrix. Our consideration for doing this is: (1) no rating may also suggest a preference; (2) this is beneficial for enhancing the model’s capability to filter noise and identify NT. With the knowledge from the source domains, the positive knowledge activation module is to alleviate NT after the knowledge transfer (CH2). Although we have aligned the feature space in the previous module, the Gaussian noise that has been fed to the target domain with source domain embedding matrices leads to potential NT. How to utilize the transferred knowledge remains a great challenge.
@@
To alleviate NT, common approaches are to generate domain attention by predefined domain features [13] or to control the transfer ratio of source domains by hyper-parameters [7]. These methods are only applicable to a limited number of domains and have excessive human intervention. In FedGCDR, we take an attention-based approach. First, we expand the $u$’s (Mary’s) local graph of the target domain as shown in Figure 3. For the source domain embedding matrices $\hat{\mathbb{X}}_{S_{1}},\hat{\mathbb{X}}_{S_{2}},...,\hat{\mathbb{X}}_{S_{M-1}}$, we represent them as $M-1$ virtual users. Since the virtual users constructed by source domain embeddings are actually the same person $u$, they have continuity in preference and their features (i.e., embeddings) characterize $u$’s preference from different dimensions. Inspired by social recommendation [45; 46; 47], we consider that there is a implicit social relationship between virtual users and the actual user $u$, because of the continuity in their preferences. Then, we build virtual social links between them to expand the original target domain graph. Second, by applying this expanded graph in target domain training, GAT generates the corresponding attention coefficient for the virtual users, which can be understood as domain attentions. Leveraging the domain attention coefficients, the target domain can focus on domains that transfer positive knowledge and we can finally give $f_{T}(\cdot)$:
@@
Through the graph expansion, we achieve: (1) dynamic domain attentions focusing on positive source domain knowledge to alleviate NT; (2) domain attentions are generated by GAT to avoid human interventions such as hyper-parameter tuning or feature engineering.
@@
After target domain training with the expanded graph, the target domain GAT model assimilated knowledge from the source domains. However, NT may still be inevitable, and negative knowledge may have accumulated in the target domain. An example is that the Gaussian noise matrices from the source domains where the user has no rating data. On the basis of this consideration, we adopt an additional fine-tuning stage: First, we freeze the message propagation layer of GAT to isolate the influence of source domains. This is because the Gaussian noise still floods through the transfer process. Second, we directly train the well-informed embeddings generated by the target domain GAT. These steps adapt the learned external knowledge for predicting the target domain ratings. In this process, we use the loss of prediction in Equation (11) as the object function:

### v2

@@ 1 Introduction
In this paper, we focus on a more generic scenario of Broader-Source Cross-Domain Recommendation (BS-CDR), which integrates knowledge from more than two source domains while preserving privacy. Given the diverse nature of user preferences, it is essential to gain a more holistic understanding of user interests by incorporating user behaviors from diversified domains  [10; 11]. For example, in Figure 1(a), a user who enjoys certain types of books might also enjoy movies, toys, and games in similar genres. However, incorporating more domains while preserving privacy poses challenges to counteract negative transfer (NT), which is a phenomenon of transferring knowledge from a source domain that negatively impacts the recommender model performance in the target domain [12]. Suppose the Books domain in Figure 1(a) is the target domain. The Clothing domain is causing NT, because of the domain discrepancy. While the Music domain is supposed to transfer positive knowledge, it might also lead to NT because of lossy privacy-preserving techniques applied to broader source domains. As a result, the influx of negative knowledge accumulated from source domains will poison the model performance of the target domain in BS-CDR scenarios.
To mitigate the NT issue, attention mechanisms have been widely leveraged, either in an explicit (e.g., determine domain attentions by predefined domain features [13; 14]) or implicit manner (e.g., employ hyper-parameters [7; 15]). Several other studies [16; 17; 18], ensure positive transfer by passing only domain-shared features. However, existing methods cannot be directly applied to BS-CDR due to two major challenges. First, inadequate privacy preservation (CH1). Both intra-domain and inter-domain privacy must be carefully considered in BS-CDR. As depicted in Figure 1(a), BS-CDR relies on extensive knowledge transfer, risking simultaneous privacy leakages across broader source domains (inter-domain privacy) [9; 19; 20; 21]. Additionally, concerns over centralized data storage may prevent users from sharing sensitive rating data (intra-domain privacy). Second, accumulative negative transfer (CH2). Adjusting attention-related hyper-parameters for a large number of source domains in BS-CDR scenarios is extremely difficult, as well as predefined or domain-shared features cannot accommodate complex domain diversities. In addition, the use of various lossy privacy-preserving techniques can further degrade the quality of transferred knowledge, complicating the achievement of positive transfer. Consequently, the impact of NT can inevitably intensify with an increasing number of source domains [2] and the performance of CDR models can decline to levels lower than those of single-domain models, as shown in Figure 1(b).
@@
Figure 1: (a) In order to obtain accurate recommendations in the Books domain, we aim to exploit user preferences (i.e., knowledge of external domains should be fully utilized, e.g. Movie, Toys, and Games domains). However, with the influence of lossy privacy-preserving techniques, the results of the transfer could be negative (e.g., the Music domain with low-quality data). (b) There is a diminishing marginal effect on the growth rate of the model performance with pure positive knowledge, while NT accumulates with an increasing number of source domains. Consequently, the performance of existing methods declines and is worse than that of a single domain model.
@@
  • We propose two key model, i.e., the positive knowledge transfer module and the positive knowledge activation module. The first transfer module ensure privacy and positive knowledge flows via privacy-preserving knowledge extraction and feature mapping. The second activation module filter harmful information via graph expansion, target domain training and target model fine-tuning;
  • We conduct extensive experiments on 16 domains of the Amazon datasets that confirm the effectiveness of FedGCDR in terms of recommendation accuracy.
@@ 2.2 Federated recommendation
Recently, FL [28; 29; 30; 31; 32] has been widely adopted to tackle the privacy issue in recommender system. Chai et al. [33] adopts FL to classic matrix factorization algorithm and utilize homomorphic encryption to avoid the potential threat of privacy disclosure. Later, Wu et al. [34] explores the application of federated graph neural networks (GNN) models to improve the recommendation quality and ensure user privacy. To utilize sensitive social information, Liu et al. [8] adopts local differential privacy (LDP) and negative sampling. More recent studies use VFL to protect company’s privacy in recommender system. Mai et al. [35] utilizes random projection and ternary quantization to ensure privacy preservation in VFL. In CDR, Chen et al. [9] designs a dual-target VFL CDR model with orthogonal mapping matrix and LDP for organizations’ privacy preservation. Liu et al. [36] designs a graph convolutional networks (GCN)-based federated framework to learn user preference distributions for more accurate recommendations. To ensure user privacy in CDR, Liu et al. [6] utilizes a VAE-based federated model to mine user preference with data stored locally. Wu et al. [7] designs a personal module and a transfer module to provide personalized recommendation while preserving user privacy. These existing works, especially federated CDR frameworks, consider only one type of privacy (intra- or inter-domain). We aim to provide both intra-domain and inter-domain privacy.
@@ 3.2.1 Overview
The overall framework of FedGCDR is shown in Figure 2. FedGCDR follows a Horizontal-Vertical-Horizontal (HVH) pipeline and its two horizontal FL stages ensure the intra-domain privacy privacy. Our two key modules focus on the vertical stage and the second horizontal stage: (1) The positive knowledge transfer module preserves the inter-domain privacy by DP and alleviates NT by feature mapping. (2) The positive knowledge activation module filters out potential harmful or conflicting knowledge from the source domains. Specifically, we expand the local graph of the target domain by virtual social links, such that the target domain graph attention network (GAT) model could generate reliable domain attention based on the expanded graph. After target domain GAT model training, we further mitigate NT by adopting a fine-tuning stage.
@@
The HVH pipeline contains three stages with switching federated settings. The first horizontal stage refers to the source domain training in which source domain servers individually interacts with its domain users (clients). The private rating information is stored within each client, while the clients exchange model and gradients to train a domain-specific global model. The next two stages correspond to our two key modules (vertical positive knowledge transfer module and horizontal positive knowledge activation module), which we will cover in detail in the following subsections. It’s important to note that the vertical positive knowledge transfer module is completely computed in each client’s user space (their personal devices), thus reducing communication overheads. This is because the needed source domain knowledge can be extracted from local source models on each client which are distributed during the first horizontal source domain training stage.
Following the HVH pipeline, we achieve: (1) Privacy enhancement. The two horizontal stages can provide intra-domain privacy preservation, while we further ensure inter-domain privacy by applying DP to the vertical stage. In the mean time, servers are not involved in the knowledge transfer process (i.e., the positive knowledge transfer module), making them unaware of user interactions in other source domains. (2) Communication efficiency. Cross-domain knowledge transfer does not require additional communication overhead.
@@
We adopt a GAT-based [37; 38; 39] federated framework as the underlying model for our intra-domain recommender system. The horizontal paradigm avoids centralized storage of user ratings to ensure intra-domain privacy (CH1). In the initial step, each user and item is offered an ID embedding of size $d$, denoted by $\mathbb{e}^{0}_{u}$, $\mathbb{e}^{0}_{v}\in\mathbb{R}^{d}$ respectively. The embedding is passed through $L$ message propagation layers [40; 41; 42]. For the $l$-th layer:
@@
In each source domain, the domain server and corresponding users collaboratively train a GAT-based federated model. The training process follows the horizontal federated learning (HFL) paradigm in which only the model and gradients are exchanged considering intra-domain privacy. We will not detail the horizontal federation model (e.g., further privacy guarantee and more high-order information) as it is a well established FL model and not our novel contribution. This model can be replaced by other GAT-based FL models [34; 44] as well.
@@ 3.2.2 Positive knowledge transfer module
After the source domain training, we obtain a series of source models in individual client’s user space. Our positive knowledge transfer module then prepares positive knowledge to be transferred from each source domains ${\mathcal{D}}^{S}$ to the target domain $\mathcal{D}^{T}$, while protecting inter-domain privacy (CH1). Specifically, suppose a individual user (client) $u$ and a source domain ${\mathcal{D}}^{S_{i}}$, we transfer the user $u$’s embedding matrix $\mathbb{X}_{S_{i}}\in\mathbb{R}^{L\times d}$ Take the row $l$ of the matrix (i.e., $\mathbb{x}^{l}_{S_{i}}$) as an example, it is the user $u$’s embedding output by the $l$-th message propagation layer ($e_{u}^{l}$). In an ideal scenario (i.e., we transfer totally positive knowledge without taking inter-domain privacy into account) [6], embedding matrices from different source domains can be directly used to enhance target domain local training in client $u$. By utilizing the source domain embeddings, $u$’s final target domain embedding $\mathbb{e}_{T}^{l}$ of layer $l$ is:
@@
where $f_{T}(\cdot)$ is the function that the target domain aggregates the knowledge of the source domains and we will give its final expression in Subsection 3.2.3. In this process, the transfer of knowledge between domains takes place entirely in the user $u$’s local space. Such a fully localized mode of knowledge transfer avoiding the additional communication overhead and potential privacy issues [6]. However, this direct embeddings transfer does not meet the privacy and NT constrains in BS-CDR scenarios.
@@
In existing CDR frameworks, the user or item embedding was shared as knowledge [9; 15; 6], which neglects inter-domain privacy. In a GNN-based approach, such direct transfers are subject to privacy attacks. Each message propagation layer can be viewed as a function with user and item embeddings as input. An attacker can easily obtain the user’s private rating matrix based on these embeddings. We apply DP to the source domain embeddings $\mathbb{x}_{S_{i}}$ [22; 45] to safeguard inter-domain privacy.
@@
User features could represent personal preferences and are influenced by domain features. The discrepancy of domains leads to the heterogeneity of feature space between domains which means that source domain embeddings cannot be utilized directly by the target domain. Man et al. [15] show that there exists an underlying mapping relationship between the latent user matrix of different domains, which can be captured by a mapping function. In order to alleviate NT, we adopt a series of MLP to explore mapping functions for each source domain. Adding Gaussian noise and feature mapping, Equation (5) becomes:
@@ 3.2.3 Positive knowledge activation module
After the aforementioned operations, the target domain obtains a list of source domain matrices $\hat{\mathbb{X}}_{S_{1}},\hat{\mathbb{X}}_{S_{2}},...,\hat{\mathbb{X}}_{S_{M-1}}$. The row of the matrices represent $MLP_{i}(\hat{\mathbb{x}}_{S_{i}}^{l})$. It is worth noting that for source domains where a user has no rating, $\hat{\mathbb{X}}_{S_{i}}$ is a Gaussian noise matrix and our motivation is: (1) no rating may also suggest a preference; (2) this is beneficial for enhancing the model’s capability to filter noise and identify NT. With the knowledge from the source domains, the positive knowledge activation module is to alleviate NT after the knowledge transfer (CH2). Although we have aligned the feature space in the previous module, the Gaussian noise that has been fed to the target domain with source domain embedding matrices leads to potential NT. How to utilize the transferred knowledge remains a great challenge.
@@
To alleviate NT, common approaches are to generate domain attention by predefined domain features [13; 46; 47] or to control the transfer ratio of source domains by hyper-parameters [7]. These methods are only applicable to a limited number of domains and have excessive human intervention. In FedGCDR, we take an attention-based approach. First, we expand $u$’s (Mary’s) local graph of the target domain as shown in Figure 3. For the source domain embedding matrices $\hat{\mathbb{X}}_{S_{1}},\hat{\mathbb{X}}_{S_{2}},...,\hat{\mathbb{X}}_{S_{M-1}}$, we represent them as $M-1$ virtual users.Since the virtual users constructed from source domain embeddings represent the same individual $u$, they share correlated preferences, with their features (i.e., embeddings) characterizing $u$’s preferences. Inspired by social recommendation [48; 49; 50], we consider that there is a implicit social relationship between virtual users and the actual user $u$, because of the correlation in their preferences. Then, we build virtual social links between them to expand the original target domain graph. Second, by incorporating this expanded graph into target domain training, the GAT model generates corresponding attention coefficients for the virtual users, which can be interpreted as domain-specific attentions. Leveraging the domain attention coefficients, the target domain can focus on domains that transfer positive knowledge and we can finally give $f_{T}(\cdot)$:
@@
Through the graph expansion, we achieve: (1) dynamic domain attentions that focus on positive source domain knowledge to alleviate NT; (2) attention generation by GAT, eliminating the need for additional interventions such as hyper-parameter tuning or feature engineering.
@@
After target domain training with the expanded graph, the target domain GAT model assimilates knowledge from the source domains. However, NT may still be unavoidable, potentially leading to the accumulation of negative knowledge in the target domain. An example of this is the Gaussian noise matrices transferred from source domains where the user has no interactions. On the basis of this consideration, we adopt an additional fine-tuning stage: First, we freeze the message propagation layer of GAT to isolate the influence of source domains preventing Gaussian noise from permeating through the transfer process. Second, we directly train the well-informed embeddings generated by the target domain GAT. These steps adapt the learned external knowledge for predicting the target domain ratings. In this process, we use the loss of prediction in Equation (11) as the object function: